OpenStudy (anonymous):

Can someone please help me!? Giving medal to the best answer. Please, please, please. It's the only one I don't know. Lines p and q are parallel. Angles one and four are supplementary angles. If the measure of angle one equals the quantity of four times x degrees and the measure of angle four equals the quantity of four times x plus twenty degrees, what is the measure of angle four? Two parallel lines, labeled p and q, are cut by a transversal. twenty degrees eighty degrees one hundred degrees one hundred twenty degrees

OpenStudy (anonymous):

Here is the Picture.

OpenStudy (bibby):

Start by converting the words into equations m<1 = 4x m<4 = 4x +20 You know they're supplementary which means they (both of the angles) add up to 180

OpenStudy (anonymous):

yes, but how do I get that? @bibby

OpenStudy (bibby):

The words "Angles one and four are supplementary angles" mean that m<1 + m<4 = 180. You can just replace the angle measurements with what I wrote out

OpenStudy (anonymous):

Okay, but how do I find x? @bibby

OpenStudy (bibby):

Can you solve an equation? m<1 = 4x m<4 = 4x +20 m<1 + m<4 = 180 4x + 4x + 20 = 180

OpenStudy (anonymous):

Please help me. I don't get it. When I do the equation, I get 10... @bibby

OpenStudy (bibby):

Well first things first. We combine like terms and get 8x +20 = 180 Subtract 20 and get 8x = 160 x = ?

OpenStudy (anonymous):

20! Omg thank you so so so much!!! <3 @bibby

OpenStudy (bibby):

That's not the answer though. Remember what the question is. What is the measurement of angle 4?

OpenStudy (anonymous):

what.

OpenStudy (anonymous):

oh. the answer is 80.

OpenStudy (bibby):

You didn't answer the question yet. You solved for what x is. You still have to plug it back in. m<4 = 4x +20 80? How'd ya get that

OpenStudy (anonymous):

the answer is 100... whoops. @bibby

OpenStudy (bibby):

a+, gold stars, whoop whoop. Great job

OpenStudy (anonymous):

omg I can't thank you enough haha thank you.@bibby

OpenStudy (anonymous):

@bibby ^^